can you please help me with the follwing Quadratic Questions.
2x^-12x+17
giving the answer in surd form.
Minimum value of 2x^+8x+7
Thank you

Answer by stanbon(75887)

Quadratic Questions.
2x^-12x+17
Use the Quadratic Formula:
x = [12 +- sqrt(144 - 4*2*-17)]/(4
---
x = [12 +- sqrt(180)]/4
---
x = [12 +- 6sqrt(5)]/4
---
x = 3 +- (3/2)sqrt(5)
==================================
Minimum value of 2x^+8x+7
---
Min occurs when x = -b/2a - -8/4 = -2
----
f(-2) = 2(4)+8(-2)+7 = -1
---
Min at (-2,-1)
===============
Cheers,
Stan H.